OpenStudy (anonymous):

Quadratic equations, which are expressed in the form of ax^2+bx+c=0, where a does not equal 0, may have how many solutions…explain

OpenStudy (campbell_st):

the quadratic will have 2 solutions. the solutions may be unequal or equal Unequal solutions \[x^2 + 4x + 3 = 0.....(x + 3)(x +1) = 0\] so the solutions are -3 and -1 equal solutions \[x^2 + 4x + 4 = 0.... (x +2)(x+2) =0....... (x+2)^2 = 0\] so there is a repeated equal solution of -2 and -2